Derek plans to retire on his 65th birthday. However, he plans to work part-time until he turns 73.00. During these years of part-time work, he will neither make deposits to nor take withdrawals from his retirement account. Exactly one year after he fully retires on the day he turns 73.0, he will begin to make annual withdrawals of $100,736.00 from his retirement account until he turns 92.00. He he will make contributions to his retirement account from his 26th birthday to his 65th birthday. To reach his goal, what must the contributions be? Assume a 5.00% interest rate. SHOW FINANCE CALCULATOR STROKES
Answer #1

Using financial calcualtor

N=19

I/Y=5%

PMT=-100736

FV=0

CPT PV=1217426.88

NOW AGAIN

N=40

I/Y=5%

FV=-1217426.88/1.05^8

CPT PMT=5444.86